May 1 18 palu al malurily. Accepted a 120-day, 13% note from Licata Company in exchange for its account receivable of $4,800.1. Prepare the journal entries to record the receivable transactions during 2019 and the necessary adjusting entry on DecembeThank you for your help and please show your work.
Answer #1
Journal Entry
Date Account Title Debit Credit
Jun-02 Cash 11791.7
Loss from sale of receivables 1550
Note Receivable 11700
Interest Income 91.7
Recourse Liability 1550
Calculation of Interest Income :
1. Interest on Licata Notes = 4800$*13%*(33/360) = 57.2$
2. Interest from Eagle Notes = 6900$*12%*(15/360) = 34.5$
Total Interest Income = 57.2+34.5 = 91.7$